#x2 + y2 - 16x + 14y + 109 = 0 is the equation of a circle with center (h, k) and radius r for: h = and k= Jo and TE
Answer #1

Solhx x++y2-16x +14y +109 = 0 22-162 +64 - 64 +y2+144 + 49 -49 tag= T > (x–8)? +(4+2)2 - 113 + 109 = 0 (x-8)2 +(4+7)2 - 4 =0